Lepton + Jets Search Philosophy

Decay mode with the greatest sensitivity is “lepton+jets”

• About 30% of t t events decay this way

• Look for one high PT lepton

• Require E̸T > 25 GeV

Problems come from physical backgrounds

• Backgrounds from “fake” leptons, etc. are manageable

• “QCD W + jet” production potentially overwhelmes top signal

• Have to find additional rejection power 1. Place more stringent kinematical requirements

2. Require presence of b quark jet – Can use semileptonic b quark decays

– Can also use large cτ of b decays ∗ This capability is unique to CDF

• Have to maintain good top efficiency

W+Jet Selection

Selection of W candidates straightforward

• Require high PT isolated lepton candidate

• Require significant E̸T

• Require three or more high PT jets

Both experiments have large samples

• CDF has 11,900 eνe and 7,000 μνμ candidates • D0 / has corresponding numbers

W + Jet Sample

Require ≥ 3 or 4 jets (from other W and b quarks)

• ET > 15 GeV and |η| < 2.0

• CDF finds 52 W+≥ 3 jet events

• D0 / finds 24 W+≥ 4 jet events

Most background sources are small:

Background  Number of Events (CDF)
Fake leptons and bb̄  5.2 ± 2.6
W W, W Z and ZZ  2.6 ± 1.2
Z → ττ, ee, μμ  4.4 ± 0.9
Total  12.2 ± 3.1

QCD W+jets production is largest expected background

Mtop (GeV/c2) Number of Events (CDF)
120  39 ± 6
140  22 ± 3
160  12 ± 2
180  6.6 ± 0.7
W+jets  50{+16 \atop −8}

Event Kinematics

Can use kinematic features to separate top from QCD background

• t t events are more spherical (acoplanarity)

Mij = ∑_{jets} \frac{pipj}{p2}

A = \frac{3}{2} Q1 where Q1 is smallest eigenvalue of M

• t t events have more energetic partons (HT)

HT = ∑_{jets}ET

D0 / has applied both techniques to W+≥ 4 jets

• Require A > 0.05 and HT > 140 GeV

• Estimate W+jet background by scaling jet multiplicities • Observe 8 events and expect 3.9 ± 1.6 from background

More on Event Kinematics

CDF uses a somewhat different approach

• Use expected distribution of jets – Jet 2 and 3 provide good signal-background separation – Form relative likelihood

log (Lrel) ≡ log(\frac{Ltop (E2 T , E3 T)}{LQCD (E2 T , E3 T)} )

• Compute using VECBOS (QCD) and Pythia (t t) MC

This technique becomes MC dependent

• Have to study effect of systematics on Lrel

• Calibrate technique on control samples – “background-enriched” W+jet sample

– W+2 jet sample where little top is expected

Results of Event Structure Analysis

Analysis indicates that W + ≥ 3 jet data is mixture of top + background

• Probability that this is a background fluctuation?

– Ask how often background would ∗ give at least 8 events with log (Lrel) > 0 ∗ when observe 14 total

• Result is 0.8%

B Tagging Techniques

Use two properties of b quarks to identify them

1. About 20% of b’s decay via

b → e−νeX and b → μ−νμX

• Look for jets with “soft” leptons in them

• Call these soft-lepton tags (SLT’s)

• CDF expects about 15% of t t events to have at least one SLT tag

• D0 / only uses soft muons

– has higher efficiency for “cascade” decays b → cX → μνμX′

– Expects to be able to tag about 20% of t t events 2. “Long” b lifetime – cτ ∼ 450 microns

• CDF uses Silicon Vertex Detector to identify jets with a secondary vertex

• We call these “SVX” tags

• Expect 24% of t t events to have at least one SVX tag

Idea is to just “count” number of tagged events

Soft Lepton Tagging

CDF identifyies muon or electron from b hadron decay

• Require lepton PT > 2 GeV/c

• Require electron to satisfy standard cuts – Use dEdx measurement in CTC to reject pions

– Require muon candidate to have good quality muon stub

D0 / searches for muons with P_{T} > 4 GeV/c

• Lepton threshold does not effect acceptance significantly

CDF has studied SLT tagging using

• photon conversions for electrons (γ → e+ e−)

• J/ψ → μ+ μ− decays

• generic jets to obtain fake rates

CDF efficiency and fake rate

• Efficiency is 0.152 ± 0.023 for Mtop = 160 GeV/c2

• “Fake” tagging rate is ∼ 0.005 − 0.010 tags/track

SVX B Tags

SVX tag requirements are

• at least 2 tracks with impact parameter d > 3σ

• track PT > 2 GeV/c

• the tracks are consistent with coming from a common vertex

Used control samples to measure the tagging efficiency and fake rate

• inclusive electrons enriched in b → e+ νeX decays

• B → J/ψX, J/ψK±, J/ψK∗ decays

• “generic” jet sample

b Quark Tagging in W+Jet Events

Applied these tagging techniques to W+jet events

• The W + 1 and W + 2 jet samples are ultimate control – Don’t expect significant top contribution

– Theory predicts low b hadron content

Sample  SVX bkg  SVX tags  SLT bkg  SLT tags
W+1 jet  15.8 ± 2.1  8  34.1 ± 3.3  33
W+2 jet  6.3 ± 0.8  8  9.9 ± 1.0  12
W+≥ 3 jet  2.30 ± 0.29  6  3.1 ± 0.3  7

CDF observes 13 tags in 10 events

• 3 events are double-tagged

• Expected to see 5.4 ± 0.4 background tags

D0 / observes 8 tagged events

• Expect 2.3 ± 0.4— background events

• Expect about 3 events if Mtop∼ 160 GeV/c2

CDF Summary of Jet Counting Experiment

We combine the SLT and SVX tagged data

• Compare with background estimates

• Observe 13 tags when we expect 5.4 ± 0.4 background

D0 / Counting Experiment

D0 / combines dilepton, kinematic and μ-tagged events

• Observe 8 μ-tagged events on background of 2.3 ± 0.4

• Observe 8 events on background of 3.9 ± 1.6 from A − HT • 2 dilepton events (expect 2.0 ± 0.2 background)

Gives a total of 18 events on background of 8.2 ± 1.7

• Probability of background fluctuation is 0.7%

• Not yet significant enough on its own

Expectations from top are

Mtop (GeV/c2) Number of Events (D0 / )
140  18.0 ± 2.0
160  11.8 ± 1.7
180  6.8 ± 1.0
200  4.0 ± 0.6

Features of CDF Tagged Samples

Find that 3 events are “double tagged”

• 2 of these 3 have 2 different jets tagged

• Probability of 2 fake tags is very low

• Expect only 1.3 double-tagged events from background

One dilepton event has jet tagged as an SLT and SVX

• Expect only 0.13 events from purely background

• Expect 0.7 double tags if both events are t t

Number of “double-tags” larger than expected from t t

• Have number of tagged-events fluctuated low?

• Number of double-tags fluctuated high?

Statistical Significance of CDF Sample

Observe 12 events in total

• 2 dilepton events when we expect 0.56{+0.25 \atop −0.13}

• 13 tags when we expect 5.4 ± 0.4

• 3 double-tagged events

Evaluated the probability of this assuming no top

• Estimate the number of tags from the background sources • Simulate our experiment assuming no top

• Count how often we would see at least the number of tags

Sample  Probability
6 SVX tags  0.032
7 SLT tags  0.038
13 tags + 2 Dilepton events  0.0026

Kinematic Features of Events

Examined the kinematics of these events

• Are there other features of data that indicate top?

• Good example are energies of second and third jet

• Need to trust MC calculations for W+jet and t t production

Kinematics suggest that of the 52 W + ≥ 3 jet events

• ∼ 20 events from t t production

• remainder mostly QCD W + jets

• Confirmed by looking at ET 2 versus ET 3 for tagged sample

The event-structure analysis provides quanitative confirmation

• Best described as combination of t t and QCD

Cross Section for Top Production

Can use our data to estimate top quark cross section

• Subtract the expected backgrounds – 1.6 ± 0.7 and 1.5 ± 0.7 SVX and SLT background tags

• Perform a maximum likelihood fit to estimate σt t – Note that counting tags gives the largest possible σ – Reanalysis of data counting events reduces it by about

one standard deviation

Measurement of Top Mass

We use event kinematics to estimate top quark mass

• Require events to have a fourth jet – Require ET > 8 and |η| < 2.4 for 4rth jet – High efficiency (∼ 85% for t t)

• Combine all possible jet pairs to form W → qq̄ – Choose one with best χ2

– Require χ2 < 10

• Use b-tag information to identify b quark jet

• Fit events requiring kinematic constraints

This selection leaves us with 7 tagged events

• Estimate background to be 1.4{+2.0 \atop −1.1} events – Primarily QCD W+jet

Procedure is dependent on theory predictions

Result of Mass Fits

The fitted mass for the 7 events is consistent with t t Fit the mass distribution to obtain

M_{top} = 174 ± 10_{−12}^{+13} GeV/c^{2}

Systematic error is dominated by uncertainties in

• Gluon radiation in t t production (4.4 GeV/c2)

• Different background shapes (∼ 5 GeV/c2) – Not clear what shape the non-W background has

Physics Comments

Perhaps the most interesting immediate result is how heavy top appears to be

• In good agreement with indirect measurements

• Consistent with MW measurements and SM

• Cross section is somewhat high

Other Possible Hypotheses

We have considered a number of alternative hypothese

1. Are we seeing a b′?

• Cross section would be too large

2. Are we seeing the Higgs boson?

• Only expect to see ∼ 0.2 events at best

3. Could be something VERY exotic (eg. H+ ) • Will simply need more data

“Best” interpretation is that we are seeing SM top

Now the Fine Print....

Statistically marginal signal

• probability of background fluctuation 0.26%

• Really would like to have ×2 more data

Two consistency checks raise some concerns

1. Predicted rate of QCD W+4 jet events too large

Process  Observed Events
Predicted QCD W+4 jet rate  8 ± 3{+6 \atop −2}
Estimated rate from t t  11.6{+5.6 \atop −4.5}
Observed W+4 jet rate  9

2. Observe an apparent excess of Z + 4 jet events

• Observe 3 Z∘ + 3 jet events and 2 Z∘ + 4 jet events • Both Z∘ + 4 jet events have b tags!

– Expect 0.64 ± 0.06 background tags

Both observations are statistics-limited

• Cannot rule out statistical fluctuation

• But cannot rationalise these two observations away either

Future Prospects

Tevatron began running again in January 1994

• Run IB will continue through end of 1995

• Maximum luminosity of 1.7 × 1031 cm−2s−1 attained • Goal is for CDF/D0 / to collect > 100 pb−1

This will allow us to confirm the “top” signal

• Both CDF and D0 / will be giving updates soon – CDF now has data sample of ∼ 65 pb−1

– Working to repeat and extend Run 1A analysis

CDF has made many improvements to analysis

• New rad-hard SVX detector and improvements in tracking – Gives ×1.3 in SVX tagging efficiency

• Enhanced SVX tagging algorithm (SECVTX)

– Factor of 1.25 improvement

Prospects for Run 1A + Run 1B

Expect statistical uncertainty on M_{top} to scale as 1/√N

• Have performed MC “toy” studies to evaluate this

• Expect statistical σM ∼ 6 GeV/c2

• Systematic σM hardest to control

– Calibrate gluon radiation using other processes

∗ Study of gamma+ jet events allow us to calibrate

energy flow out of jets

– Sensitivity to background shape also should go as 1/√N ∗ Will have enough statistics to measure shape of non-

W background

• Reasonable to expect systematic uncertainty to drop as 1/√N

Can begin to estimate couplings

• Will have of order 10 dilepton events

• Expect of order 50 lepton + jet events

• Might expect ∼ 30% measurement of

\frac{BR(top → leptons )}{BR(top → hadrons )}  (1)

Prospects for Run II

Parameters for Run II coming into focus

• Expect maximum luminosity of 2 × 1032 cm−2s−1

• Total dataset of 1000 pb−1

• Run from 1999 to 2003 (?)

D0 / and CDF upgrades well along for Run II

• For CDF, – Upgraded Plug/Forward calorimeter system – More complete SVX coverage (×2)

– New trigger and electronics

– New central tracker…

• For D0 / , – Solenoid and new central tracker

– SVX detector (at last!)

– New electronics, trigger and DAQ system

• Overall upgrade project of order $100 M

Run II Top Physics

Top physics would enter a new era

• Expect of order 300 fully reconstructed t t decays

• How well can we measure the mass?

– Naive estimate is σstat ∼ 3 GeV/c2

– Systematic uncertainty should scale similarily

– Should realise better than σM ∼ 5 GeV/c2

• Uncertainties in semileptonic branching ratios should be better than 20%

– Hard to predict without better estimate of cross section – But look at bottom SL BR uncertainty for reality check

Conclusions

CDF and D0 / see signals consistent with top

• Most channels show an excess over background

• CDF kinematic analyses provide additional confirmation • But still at best only ∼ 3σ effect

• Probability of background fluctuation:

– CDF counting experiment significance 0.25% – D0 / counting experiment significance 0.7% – CDF event-structure analysis 0.8%

Best estimate of top quark mass is M_{top} = 174 ± 10_{−12}^{+13} GeV/c^{2}

However:

• Need one experiment that gives ≥ 5σ effect

• Published CDF results have some contraindications – Theoretical W +jet rate appears too large – Details of Z+jet production raise questions

Stayed tuned as more results will be coming shortly